Home‎ > ‎Reports‎ > ‎Crew099‎ > ‎

day07


Crew 99 Log Book for February 4, 2011

Photos of the day





Commander's Check-In Report

Donna Viola Reporting


Crew Physical Status: Good

Time Departed/Returned from EVA: EVA #10 9:30-11:15; EVA #11 15:35-17:00

Brief Narrative of Field Mission Results: marking road junction
waypoints, EVA planner test

EVA Data/Interpretations: as per EVA reports

Engineering/Hab Maintenance: as per engineering report; outside water
tank now full, yay!

Report Transmission Schedule: Engineering, EVA #10, EVA #11, Chef's,
Commander's. By 8:30pm

Plans for Tomorrow: to be decided; will probably do another EVA to
continue collecting coordinates of road junctions

Inventory: n/a

Support Requested:

Miscellaneous:



Commander's Report

Donna Viola Reporting


The crew seems to have settled into quite the routine out here - though living on Mars ought to be anything BUT routine! This morning's EVA was mostly to find some of the coordinates of several road junctions to make travel easier both for ourselves and for future crews, with Heidi, Kevin, and Max setting out on the ATVs for a two hour adventure - staying, of course, on actual roads.

In the afternoon, Andie taught Kevin how to use the EVA planner; then they went to explore Area 42. Once they neared the Hab, Heidi joined them for a photo op. Sukrit once again spent most of the day preparing to use the observatory.

Since tonight looks like the last clear night for a few days, the astronomers are planning on heading outside after dinner to try to align the telescope, and take some pretty pictures if there's enough time.

Also, Max has continued prepping for his balloon launch. Currently, it's looking like the best night to launch is going to be next Thursday.

Right now, today's chefs, Andie and Heidi, are busily preparing stir-fry for dinner. It smells delicious - but, sadly, it won't be ready for a little while yet.


Engineering Report

Kevin Newman Reporting

Generator/Electricity:


· Kitty (Diesel Generator) (Used? Yes/No): Yes


· Honey (Gas Generator) (Used? Yes/No): No


· Battery Charge Level:


o 10.59 V everstart


o 12.37 V armor plate


o 13.32 V green


Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

1 generator run, all is well


Fuel Status:


· Diesel (Full/0.75/0.5/0.25/Empty):2/7


· Propane Tank (Full/0.75/0.5/0.25/Empty):77%


· Gasoline Tank (Full/0.75/0.5/0.25/Empty): 9 Gallons


Notes/Comments:


Oil:


· Kitty Oil Quantity (# of Quarts in storage): Leave it to DG


· Honey Oil Quantity (# of Quarts in storage): Leave it to DG


· ATV Oil Quantity (# of Quarts in storage): 6 quarts on the engineering work bench


· Car Oil Quantity (# of Quarts in storage):


Water Status:


· Outside Potable Water Tank Level (inches from bottom): 15


· Trailer Potable Water Tank Level (inches from bottom): FULL


· Hab Potable Water Tank Level (inches from bottom): 9


· Potable Water Meter Reading: 416708


Notes/Comments:

The trailer mysteriously disappeared at some point in the day, only to re-appear later filled with water. We have two hypotheses for these anomalies. 1) It was temporarily stolen by Jawas. 2) DG is somewhat of a ninja. Thank you DG! However the trailer potable water tank screw cap has gone missing…


GreenHab:


· Condition of plants in Tank 1: Nothing


· Condition of plants in Tank 2: Ok


· Condition of plants in Tank 3: Good


· Greenhab notes/ comments:


Observatory:


· Telescope Used? (Yes/no) Yes


Observatory Notes:


Transportation:


Hab Car (NOT RENTAL)


· Used (yes/no):No


· Oil Checked (yes/no): No


· Tire Status: Not checked


· Odometer at end of day: 123597.9


· Notes/Comments on Hab Car:


ATV 1 (Opportunity):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: Replaced


· Tire Status: Checked, Filled


· Comments:


ATV 2 (Spirit):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: Replaced


· Tire Status: Checked, Filled


· Comments:


ATV 3 (Viking I):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: None


· Tire Status: Checked


· Comments: Started briefly using pull cord for the sake of keeping the engine in good shape.


ATV 4 (Viking II):


· Used: Yes


· Oil Checked: Yes


· Fuel Consumed: Replaced


· Tire Status: Checked, Filled


· Comments:


Heating and Ventilation:


· Thermostat setting upstairs (Farenheit): 66


· Actual temperature upstairs (Farenheit): 80


· Thermostat setting downstairs (Farenheit): 65


· Actual temperature downstairs (Farenheit): 67


Computers/Networking Infrastructure:


Notes/comments on internet/computers?: All good


General Engineering

1. Notes/Comments:


About 300ft north-east of the Hab, I found several stones oddly placed in a hemisphere. I thought they would be useful for holding down the ATV covers, so I moved them to the front of the hab.


Also, are there replacement light bulbs for the 18V Dewalt portable lantern? The part number is BLC 60018X.


EVA 10 Report

Kevin Newman Reporting

Time: 09:30-11:15

Crew Members: Max Fagin, Heidi Beemer, Kevin Newman

Site Location:  Mars

 Transit Mode: ATV

Objectives: Determine coordinates of nearby road intersections for shared MDRS database.

Results: Successfully mapped tracks and intersection coordinates for Cactus Road, Sagan Street, and Radio Ridge Road.

Lessons Learned: Exploring is fun, and providing waypoints for ourselves and future groups to aid in local navigation will be very helpful.


EVA 11 Report

Andie Gilkey Reporting

Time: 15:30-17:00


Crew Members: Andie Gilkey, Kevin Newman


Site Location:  UTM: 518850 4251000

 Transit Mode: foot


Objectives: Measure how accurately the mission planner predicts astronaut energy consumption, distance traveled, and EVA time on the way to Area 42.


Results: We were able to successfully measure astronaut heart rate throughout the length of the mission, as well as distance traveled and EVA time.


Lessons Learned: There are Martian cows in the area because we saw mysterious space cow prints in the snow.


Cook's Report

Heidi Beemer Reporting

1a) Today was

-A cooking day.


1b) Was today a special day celebrated at the main meal with special

food or activity?

- Yes, today was earth’s conjunction and we celebrated with festive food, song, and dance.


2)  List the foods served at today's epic meals , giving full names of each.

- Lunch: Grilled Cheese and Bear Creek Chicken noodle soup.

        Grilled cheese was prepared from Martian made white bread and Velveeta cheese slices grilled on stove.

-Dinner: Stir Fry over Brown rice. See recipe below.


       

3) List any main meal foods not finished by the crew (leftovers stored

for later use or discarded) 

-8 servings of rice and 10 servings of rehydrated not cooked veggie salad was stored in fridge.


4)  Recipes for anything more complicated than following package

directions exactly, or rehydrating.

Ingredients

•       4 (4 ounce) boneless skinless chicken breast halves

•       3 tablespoons cornstarch

•       2 tablespoons soy sauce

•       1/2 teaspoon ground ginger

•       1/4 teaspoon garlic powder

•       3 tablespoons cooking oil, divided

•       2 cups broccoli florets

•       1 cup sliced celery (1/2 inch pieces)

•       1 cup thinly sliced carrots

•       1 small onion, cut into wedges

•       1 cup water

•       1 teaspoon chicken bouillon granules


Note: This serves 4 so was added more to fit our needs.

Tip 1: A cup of rehydrated vegetables is less than a cup of hydrated vegetables.

Tip 2: A cup of non-cooked rice is equal to three cups of cooked rice

Tip 3: Rice takes 50 mins. to cook not 10

Directions

1.      Rehydrate Chicken; place in a resealable plastic bag. Add cornstarch and toss to coat. Combine soy sauce, ginger and garlic powder; add to bag and shake well.

2.      In a large pan heat 2 tablespoons of oil; stir- fry broccoli, celery, carrots and onion for 4-5 minutes or until crisp-tender. Add water and bouillon. Add chicken to pan. Cook and stir until thickened and bubbly.

3.      Add bouillon juice to rice for chicken flavor


5)  Comments/questions for Kim and Jean, the food study investigators

- None


Journalist Report

Sukrit Ranjan Reporting

Today was a quiet day. It saw more EVA software testing and general surveys. In the Hab, further progress was made on astronomy software.


Lunch saw delicious grilled cheese, and dinner stands witness to an ongoing stir-fry experiment with - ahem - somewhat surplus amounts of rice.


Tonight is supposed to be the last cloudless night in a while, so the astronomers are going to go out and go all out. Onward!